# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 1. Introduction
An initial data set is a triple $(M,g,q)$, where $M$ is a manifold, $g$ is a Riemannian metric on $M$, and $q$ is a symmetric $(0,2)$-tensor on $M$. Denote the scalar curvature of $g$ by $R$. We define
@@
In this paper we will formulate conditions which insure that a given initial data set is of the form described in this example. The study of rigidity for initial data sets under the dominant energy condition can be traced back to Schoen and Yau’s work [SY81], in which they proved the spacetime positive energy theorem using minimal surface techniques in dimension 3. Eichmair, Huang, Lee, and Schoen [EHLS] later extended this work by proving a more general spacetime positive mass theorem up to dimension 7. Huang and Lee [HL] subsequently affirmed the rigidity statement of this theorem studied in [EHLS]. In a different approach, Witten [Wit] used spinors to prove the spacetime positive mass theorem, a method that generalizes to all higher dimensions for spin manifolds. Parker and Taubes [PT] gave a mathematically rigorous formulation of Witten’s argument. Chruściel and Maerten [CM] later provided a rigorous proof of the desired rigidity statement in all dimensions for spin manifolds, building on Beig and Chruściel’s proof [BC] in dimension 3. Lee’s book [Lee] offers a comprehensive exposition on the positive mass theorem.
@@
Let $K\subset\mathbb{R}^{n}$ be a compact convex domain with smooth boundary $\partial K$ and exterior normal $N\colon\partial K\to S^{n-1}$. Note that $N$ has degree $+1$. Let $f\colon K\to\mathbb{R}$ be a smooth map satisfying
@@
Let
[…]
be the graph of $f$. Then $M$ is a compact smooth spacelike hypersurface with boundary $\Sigma=\partial K\times\{0\}\subset\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n,1}$, see Figures 1 and 2.
@@
Let $I\!I$ denote the vector-valued second fundamental form of $\Sigma$, viewed as a submanifold of $\mathbb{R}^{n,1}$ of codimension $2$. Then
$I\!I(X,Y)=-\langle dN(X),Y\rangle\,N$
@@
We denote by $T$ the future-oriented normal field to $M$ in $\mathbb{R}^{n,1}$, where $T$ is normalized so that $\langle T,T\rangle=-1$. Let $g$ denote the induced Riemannian metric and let $q$ denote the second fundamental form of $M$ with respect to $T$. In other words, $q(X,Y)=\langle\bar{D}_{X}T,Y\rangle$ for all $X,Y\in TM$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$. The initial data set $(M,g,q)$ satisfies the dominant energy condition, cf. Example 1.1. At each point on $\Sigma$, we denote by $\nu$ the outward-pointing unit normal vector to $\Sigma$ in $M$. Moreover, we denote by $h$ the second fundamental form of $\Sigma$, viewed as a hypersurface in $M$, with respect to $\nu$. In other words, $h(X,Y)=\langle D_{X}\nu,Y\rangle$ for $X,Y\in T\Sigma$, where $D$ denotes the Levi-Civita connection on $(M,g)$. We will use these sign conventions for $q$ and $h$ throughout our paper.
@@
$I\!I(X,Y)=-h(X,Y)\,\nu+q(X,Y)\,T$
for all tangent vectors $X,Y\in T\Sigma$. Putting these facts together, we conclude that
$h(X,Y)\,\nu-q(X,Y)\,T=\langle dN(X),Y\rangle\,N$
for all tangent vectors $X,Y\in T\Sigma$. In particular,
$h(X,Y)=\langle dN(X),Y\rangle\,\langle N,\nu\rangle$
@@
$q(X,Y)=\langle dN(X),Y\rangle\,\langle N,T\rangle$
for all tangent vectors $X,Y\in T\Sigma$. We now use Assumption (c). Hence, at each point on $\Sigma$, we either have $dN=0$, or else $\langle N,\nu\rangle=1$ and $\langle N,T\rangle=0$. Consequently, $h(X,Y)=\langle dN(X),Y\rangle$ and $q(X,Y)=0$ for all tangent vectors $X,Y\in T\Sigma$. In particular, $H=\|dN\|_{\mathrm{tr}}$ and $\mathrm{tr}(q)-q(\nu,\nu)=0$ at each point on $\Sigma$. Here, $H$ denotes the unnormalized mean curvature, i.e. the sum of the principal curvatures, of $\Sigma$ in $M$. In other words, $H$ is equal to the trace of $h$, considered as an endomorphism field. Moreover, $\|\cdot\|_{\mathrm{tr}}$ denotes the trace norm of a linear map, i.e., the sum of all singular values.
@@
Let $(M,g,q)$ be an initial data set. Assume that $M$ is a compact, connected spin manifold, has dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that
  • $(M,g,q)$ satisfies the dominant energy condition,
  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$,
  • $N$ has positive degree.
@@
The quantity $H-|\mathrm{tr}(q)-q(\nu,\nu)|$ is related to the notion of marginally trapped surfaces in general relativity. Recall that a marginally outer trapped surface (MOTS) is a hypersurface in an initial data set $(M,g,q)$ satisfying $H+(\mathrm{tr}(q)-q(\nu,\nu))=0$; see [EHLS]*p. 89.
@@
There is also a variant of Theorem I for $n=1$. In this case, $M=[a,b]$ is a compact interval, $g=dt\otimes dt$ is the standard metric, and $q=\kappa g$ where $\kappa\colon[a,b]\to\mathbb{R}$ is a given smooth function. We can choose $N\colon\partial M=\{a,b\}\to S^{0}=\{-1,1\}$ to be bijective. Then all quantities $R$, $\mu$, $H$, $J$, $dN$, and $\mathrm{tr}(q)-q(\nu,\nu)$ vanish. Thus the assumptions of Theorem I are trivially satisfied.
The conclusion is now that there is a spacelike curve $[a,b]\to\mathbb{R}^{1,1}$, parametrized by arc-length and with curvature $\kappa$. Indeed, this can be shown directly by ODE methods. The analogue for the Euclidean plane $\mathbb{R}^{2}$ instead of $\mathbb{R}^{1,1}$ is known as the fundamental theorem of planar curve theory.
@@
Let $(M,g)$ be a compact connected Riemannian spin manifold of dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that
  • $R\geq 0$ holds on $M$,
  • $H\geq\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$,
  • $N$ has positive degree.
Then $M$ can be isometrically embedded in $\mathbb{R}^{n}$ as a compact convex domain with smooth boundary and exterior normal $N$.
In the second part of this paper, we consider initial data sets of the form $(\Omega,g,q)$, where $\Omega$ is a compact, convex polytope in $\mathbb{R}^{n}$ with non-empty interior. When $q=0$, the dominant energy condition simplifies to the non-negativity of the scalar curvature $R\geq 0$, and the boundary condition reduces to mean convexity $H\geq 0$, allowing us to compare $(\Omega,g)$ with Euclidean polytopes. Gromov [Gro14] initiated the study of scalar curvature comparison for polytopes with nonnegative scalar curvature to explore the notion of scalar curvature lower bounds in low-regularity spaces, and this problem has been extensively investigated; see [Gro14, Gro19, Gro22, Li17, Li19, WXY]. For $q=g$, the dominant energy condition translates to $R\geq-n(n-1)$. Li [Li20] established a polyhedral comparison result for metrics with scalar curvature at least $-n(n-1)$.
To explore the notion of energy density lower bounds in spacetimes with low regularity, we extend Gromov’s initial work [Gro14] to initial data sets under the dominant energy condition. In the case where $\Omega$ is a cube, Tsang [Tsang] addressed this problem in dimension 3 using the spacetime harmonic function approach.
We will employ the strategy developed by the second named author in [Brendle], adapting it to initial data sets. To achieve this, we utilize a boundary value problem for Dirac operators on smooth domains, which have been extensively studied thanks to the work of Hörmander [Hormander], and the first named author and Ballmann [BB, BB13]. We write the given polytope in the form $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and $u_{i}$ are linear functions on $\mathbb{R}^{n}$. For each $i\in I$, let $N_{i}\in S^{n-1}$ and $\nu_{i}$ denote the outward pointing unit normal vectors to the half-space $\{u_{i}\leq 0\}$ with respect to the Euclidean metric and the metric $g$, respectively. We make the following assumption:
@@
Assume that $n\geq 2$ is an integer. Let $\Omega$ be a compact, convex polytope in $\mathbb{R}^{n}$ with nonempty interior. Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$. Let $q$ be a symmetric $(0,2)$-tensor, which is defined on an open set containing $\Omega$. We assume that
  • $(\Omega,g,q)$ satisfies the dominant energy condition,
  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|$ on the boundary faces of $\Omega$,
@@
Conjecture 1.7.
Under the assumptions of Theorem II, there exists a smooth map $f\colon\Omega\to\mathbb{R}$ satisfying
  • $|df|<1$ on $\Omega$,
  • $f=0$ along $\partial\Omega$,
and an isometry of $(\Omega,g)$ to
$\textrm{graph}(f)=\{(x,f(x))\colon x\in\Omega\} \subset\mathbb{R}^{n,1}$
with the induced Riemannian metric. Furthermore, under this isometry, $q$ is the second fundamental form of $\textrm{graph}(f)$ in $\mathbb{R}^{n,1}$.
@@ 2. A boundary value problem for the Dirac operator in odd dimensions
Throughout this section, we assume that $M$ is a compact, connected spin manifold of dimension $n\geq 2$ with boundary $\partial M=:\Sigma$. Let $g$ be a Riemannian metric on $M$. Suppose that $\nu$ is the unit outward normal field along $\partial M$ and that $N\colon\partial M\to S^{n-1}$ is a smooth map.
@@
Let $S_{0}$ denote the spinor module of the complex Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})$ of $\mathbb{R}^{n}$. Now both $S_{0}$ and its dual space $S_{0}^{*}$ are $m$-dimensional Hermitian vector spaces.
@@
(2.1) $\mathcal{E}|_{x}=\mathcal{S}|_{x}\otimes S_{0}^{*}=\mathrm{Hom}(S_{0},\mathcal{S}|_{x}).$
Thus, a section $s$ of $\mathcal{E}$ can be considered as a field of homomorphisms. Given an element $\sigma\in S_{0}$ we can apply $s$ at each point $x\in M$ to it and obtain a section $x\mapsto s(x)\sigma$ of $\mathcal{S}$.
Denote by $\gamma\colon TM\to\mathrm{End}(\mathcal{S})$ the Clifford multiplication of $(M,g)$ and by $\gamma_{0}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0})$ the one of $\mathbb{R}^{n}$. This gives an induced Clifford multiplication $\gamma_{0}^{*}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0}^{*})$, where $\gamma_{0}^{*}(v)$ is defined as the adjoint of $\gamma_{0}(v)$. We define
@@
Since $\gamma(\nu)^{2}=\gamma_{0}^{*}(N)^{2}=-1$ we have
$\chi^{2}=1.$
Since both $\gamma(\nu)$ and $\gamma_{0}^{*}(N)$ are skew-adjoint we get
$\chi^{*}=\chi.$
For $\xi\in T\partial M$ the linear map $\gamma(\xi)\otimes 1$ anticommutes with $\chi$ and is an isomorphism if $\xi\neq 0$. Therefore the ±$\pm 1$-eigenspaces of $\chi$ have the same dimension.
Let $\mathcal{D}$ denote the spinorial Dirac operator on $M$, acting on sections of $\mathcal{E}$. This Dirac operator is formally selfadjoint. By Corollary 7.23 in [BB], the conditions $\chi s=s$ and $\chi s=-s$ both form elliptic boundary conditions for $\mathcal{D}$. Since $\gamma(\nu)\otimes 1$ commutes with $\chi$, the boundary conditions $\chi s=s$ and $\chi s=-s$ are adjoint to each other, see [BB]*Section 7.2.
We denote by $L^{2}(M,\mathcal{E})$ the Hilbert space of square integrable sections of $\mathcal{E}$ and by $H^{1}(M,\mathcal{E})$ the Sobolev space of $L^{2}$-sections whose distributional derivatives are also square integrable. If $\mathcal{F}\subset\mathcal{E}|_{\Sigma}$ is a subbundle of $\mathcal{E}$ along the boundary, we write $H^{1}(M,\mathcal{E},\mathcal{F}):=\{u\in H^{1}(M,\mathcal{E})\mid u|_{\Sigma}\in L^{2}(\Sigma,\mathcal{F})\}$. Note that the restriction map extends uniquely to a bounded linear map $H^{1}(M,\mathcal{E})\to L^{2}(\Sigma,\mathcal{E})$ by the trace theorem.
@@
where $\mathcal{F}^{\pm}$ are the eigensubbundles of $\chi$ corresponding to the eigenvalues ±$\pm 1$. We need the following index computation, which generalizes the results in [Brendle].
@@
Let $M$ have odd dimension $n\geq 3$. Then the operator

### v4

@@ 1. Introduction
An initial data set is a triplet $(M,g,q)$, where $M$ is a manifold, $g$ is a Riemannian metric on $M$, and $q$ is a symmetric $(0,2)$-tensor on $M$. We denote the scalar curvature of $g$ by $R$. We define
@@
The study of initial data sets under the dominant energy condition can be traced back to Schoen and Yau’s work [Schoen-Yau], in which they proved the spacetime positive energy theorem using minimal surface techniques in dimension 3. Eichmair, Huang, Lee, and Schoen [Eichmair-Huang-Lee-Schoen] later extended this work by proving a more general spacetime positive mass theorem up to dimension 7. Huang and Lee [Huang-Lee] subsequently proved the rigidity statement in this theorem. In a different approach, Witten [Witten] used spinors to prove the spacetime positive mass theorem, a method that generalizes to all higher dimensions for spin manifolds. Parker and Taubes [Parker-Taubes] gave a mathematically rigorous formulation of Witten’s argument. Chruściel and Maerten [Chrusciel-Maerten] later provided a rigorous proof of the desired rigidity statement in all dimensions for spin manifolds, building on Beig and Chruściel’s proof [Beig-Chrusciel] in dimension 3. Lee’s book [Lee] offers a comprehensive exposition on the positive mass theorem.
Instead of noncompact asymptotically flat manifolds, we consider compact manifolds with boundary. The condition of asymptotic flatness is replaced by suitable boundary conditions. In the interior, we still demand the dominant energy condition. First rigidity results have been obtained in this context by Eichmair, Galloway and Mendes in [Eichmair-Galloway-Mendes] and by Glöckle in [Gloe].
@@
Let $K\subset\mathbb{R}^{n}$ be a compact convex domain with smooth boundary $\partial K$ and exterior normal $N\colon\partial K\to S^{n-1}$. Note that $N$ has degree $1$. Let $f\colon K\to\mathbb{R}$ be a smooth map satisfying the following conditions:
@@
Let $M:=\{(x,f(x))\colon x\in K\}\subset\mathbb{R}^{n,1}$ denote the graph of $f$. Then $M$ is a compact smooth spacelike hypersurface with boundary $\Sigma=\partial K\times\{0\}\subset\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n,1}$, see Figures 1 and 2.
@@
Let $\mathrm{II}$ denote the vector-valued second fundamental form of $\Sigma$, viewed as a submanifold of $\mathbb{R}^{n,1}$ of codimension $2$. Then
$\mathrm{II}(X,Y)=-\langle dN(X),Y\rangle\,N$
@@
We denote by $T$ the future-oriented normal vector field to $M$ in $\mathbb{R}^{n,1}$, where $T$ is normalized so that $\langle T,T\rangle_{\mathbb{R}^{n,1}}=-1$. Let $g$ denote the induced Riemannian metric and let $q$ denote the second fundamental form of $M$ with respect to $T$. In other words, $q(X,Y)=\langle\bar{D}_{X}T,Y\rangle_{\mathbb{R}^{n,1}}$ for all $X,Y\in TM$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$. The initial data set $(M,g,q)$ satisfies the dominant energy condition, cf. Example 1.1. At each point on $\Sigma$, we denote by $\nu$ the outward-pointing unit normal vector to $\Sigma$ in $M$. Moreover, we denote by $h$ the second fundamental form of $\Sigma$, viewed as a hypersurface in $M$. In other words, $h(X,Y)=\langle D_{X}\nu,Y\rangle_{\mathbb{R}^{n,1}}$ for $X,Y\in T\Sigma$, where $D$ denotes the Levi-Civita connection on $(M,g)$. We will use these sign conventions for $q$ and $h$ throughout our paper.
@@
$h(X,Y)\,\nu-q(X,Y)\,T=-\mathrm{II}(X,Y)=\langle dN(X),Y\rangle\,N$
for all tangent vectors $X,Y\in T\Sigma$. As $\langle\nu,\nu\rangle_{\mathbb{R}^{n,1}}=1$, $\langle T,T\rangle_{\mathbb{R}^{n,1}}=-1$, and $\langle\nu,T\rangle_{\mathbb{R}^{n,1}}=0$, it follows that
$h(X,Y)=\langle dN(X),Y\rangle\,\langle N,\nu\rangle_{\mathbb{R}^{n,1}}$
@@
$q(X,Y)=\langle dN(X),Y\rangle\,\langle N,T\rangle_{\mathbb{R}^{n,1}}$
for all tangent vectors $X,Y\in T\Sigma$. Assumption (c) implies that, at each point on $\Sigma$, we either have $dN=0$, or $\langle N,\nu\rangle_{\mathbb{R}^{n,1}}=1$ and $\langle N,T\rangle_{\mathbb{R}^{n,1}}=0$. Consequently, $h(X,Y)=\langle dN(X),Y\rangle$ and $q(X,Y)=0$ for all tangent vectors $X,Y\in T\Sigma$. In particular, we have $H=\|dN\|_{\mathrm{tr}}$ and $\mathrm{tr}(q)-q(\nu,\nu)=0$ at each point on $\Sigma$. Here, $H$ denotes the unnormalized mean curvature, i.e. the sum of the principal curvatures of $\Sigma$ in $M$. In other words, $H$ is equal to the trace of $h$, considered as an endomorphism field. Moreover, $\|\cdot\|_{\mathrm{tr}}$ denotes the trace norm of a linear map, i.e. the sum of all singular values.
@@
Let $(M,g,q)$ be an initial data set. Assume that $M$ is a compact connected spin manifold, has dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that the following conditions are satisfied:
  • $(M,g,q)$ satisfies the dominant energy condition.
  • $H\geq\|dN\|_{\mathrm{tr}}+|\mathrm{tr}(q)-q(\nu,\nu)|$ holds along $\Sigma$.
  • The map $N\colon\Sigma\to S^{n-1}$ has positive degree.
@@
The quantity $H-|\mathrm{tr}(q)-q(\nu,\nu)|$ is related to the notion of marginally trapped surfaces in general relativity. Recall that a marginally outer trapped surface (MOTS) is a hypersurface in an initial data set $(M,g,q)$ satisfying $H+(\mathrm{tr}(q)-q(\nu,\nu))=0$; see [Eichmair-Huang-Lee-Schoen]*p. 89.
@@
There is also a variant of Theorem I for $n=1$. In this case, $M=[a,b]$ is a compact interval, $g=dt\otimes dt$ is the standard metric, and $q=\kappa g$ where $\kappa\colon[a,b]\to\mathbb{R}$ is a given smooth function. We can choose $N\colon\partial M=\{a,b\}\to S^{0}=\{-1,1\}$ to be bijective. Then all quantities $R$, $\mu$, $H$, $J$, dNdN, and $\mathrm{tr}(q)-q(\nu,\nu)$ vanish. Thus the assumptions of Theorem I are trivially satisfied.
The conclusion is now that there is a spacelike curve $a,b]\to\mathbb{R}^{1,1}$, parametrized by arc-length and with curvature $\kappa$. Indeed, this can be shown directly by ODE methods. The analogue for the Euclidean plane $\mathbb{R}^{2}$ instead of $\mathbb{R}^{1,1}$ is known as the fundamental theorem of planar curve theory.
@@
Let $(M,g)$ be a compact connected Riemannian spin manifold of dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that the following conditions are satisfied:
  • $R\geq 0$ holds on $M$.
  • $H\geq\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$.
  • The map $N\colon\Sigma\to S^{n-1}$ has positive degree.
Then $M$ can be isometrically embedded in $\mathbb{R}^{n}$ as a compact convex domain with smooth boundary. Moreover, the map $N$ agrees, up to an element of $O(n)$, with the exterior normal $\partial K\to S^{n-1}$.
The proof of Theorem I uses a boundary value problem for the Dirac operator. Boundary value problems for the Dirac operator have been extensively studied thanks to the work of Hörmander [Hormander], and the first named author and Ballmann [Baer-Ballmann, Baer-Ballmann13]. In the second part of this paper, we consider initial data sets of the form $(\Omega,g,q)$, where $\Omega$ is a compact convex polytope in $\mathbb{R}^{n}$ with nonempty interior. When $q=0$, the dominant energy condition simplifies to the nonnegativity of the scalar curvature $R\geq 0$, and the boundary condition reduces to mean convexity $H\geq 0$, allowing us to compare $(\Omega,g)$ with Euclidean polytopes. Gromov [Gromov14] initiated the study of scalar curvature comparison for polytopes with nonnegative scalar curvature to explore the notion of scalar curvature lower bounds in low-regularity spaces, and this problem has been extensively investigated; see [BW23, Gromov14, Gromov19, Gromov22, Li17, Li17Err, Li19, Wang-Xie-Yu]. For $q=g$, the dominant energy condition reduces to the inequality $R\geq-n(n-1)$. Li [Li20] established a polyhedral comparison result for metrics with scalar curvature at least n(n1)-n(n-1). The case of general initial data sets was addressed by Tsang [Tsang], in the special case when $\Omega$ is a three-dimensional cube.
We will employ the strategy developed by the second named author in [Brendle], adapting it to initial data sets. We write the given polytope in the form $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and $u_{i}$ are linear functions on $\mathbb{R}^{n}$. For each $i\in I$, let $N_{i}\in S^{n-1}$ and $\nu_{i}$ denote the outward pointing unit normal vectors to the half-space $\{u_{i}\leq 0\}$ with respect to the Euclidean metric and the metric $g$, respectively. We make the following assumption:
@@
Assume that $n\geq 2$ is an integer. Let $\Omega$ be a compact convex polytope in $\mathbb{R}^{n}$ with nonempty interior. Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$. Let $q$ be a symmetric $(0,2)$-tensor, which is defined on an open set containing $\Omega$. We assume that the following conditions are satisfied:
  • $(\Omega,g,q)$ satisfies the dominant energy condition.
  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|$ on the boundary faces of $\Omega$.
@@ 2. A boundary value problem for the Dirac operator in odd dimensions
Throughout this section, we assume that $M$ is a compact connected spin manifold of dimension $n\geq 2$ with boundary $\partial M=\Sigma$. Let $g$ be a Riemannian metric on $M$. Suppose that $\nu$ is the unit outward normal field along $\partial M$ and that $N\colon\partial M\to S^{n-1}$ is a given smooth map.
@@
Let $S_{0}$ denote the spinor module of the complex Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})$ of $\mathbb{R}^{n}$. Both $S_{0}$ and its dual space $S_{0}^{*}$ are $m$-dimensional Hermitian vector spaces.
@@
$\mathcal{E}|_{x}=\mathcal{S}|_{x}\otimes S_{0}^{*}=\mathrm{Hom}(S_{0},\mathcal{S}|_{x}).$
Thus, a section $s$ of $\mathcal{E}$ can be considered as a field of homomorphisms. Given an element $\sigma\in S_{0}$ we can apply $s$ at each point $x\in M$ to $\sigma$ and obtain a section $s\sigma$ of $\mathcal{S}$.
We denote by $\gamma\colon TM\to\mathrm{End}(\mathcal{S})$ and $\gamma_{0}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0})$ the Clifford multiplication of $(M,g)$ and the one of $\mathbb{R}^{n}$, respectively. This gives an induced Clifford multiplication
$\gamma_{0}^{*}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0}^{*}),$
where $\gamma_{0}^{*}(\xi)$ is defined as the adjoint of $\gamma_{0}(\xi)$. We define
@@
Using the identity $\gamma(\nu)^{2}=\gamma_{0}^{*}(N)^{2}=-\textup{id}$, we obtain $\chi^{2}=\textup{id}$. Since both $\gamma(\nu)$ and $\gamma_{0}^{*}(N)$ are skew-adjoint, we have $\chi^{*}=\chi$.
For each $X\in T\Sigma$, the linear map $\gamma(X)\otimes\textup{id}$ anticommutes with $\chi$ and is an isomorphism if $X\neq 0$. Therefore the $\pm 1$-eigenspaces of $\chi$ have the same dimension.
Let $\mathcal{D}$ denote the Dirac operator on $M$, acting on sections of $\mathcal{E}$. This Dirac operator is formally selfadjoint. By [Baer-Ballmann]*Corollary 7.23, the conditions $\chi s=s$ and $\chi s=-s$ both form elliptic boundary conditions for $\mathcal{D}$. Since $\gamma(\nu)\otimes\textup{id}$ commutes with $\chi$, the boundary conditions $\chi s=s$ and $\chi s=-s$ are adjoint to each other, see [Baer-Ballmann]*Section 7.2.
We denote by $L^{2}(M,\mathcal{E})$ the Hilbert space of square integrable sections of $\mathcal{E}$ and by $H^{1}(M,\mathcal{E})$ the Sobolev space of $L^{2}$-sections whose distributional derivatives are also square integrable. If $\mathcal{F}\subset\mathcal{E}|_{\Sigma}$ is a subbundle of $\mathcal{E}$ along the boundary, we write
$H^{1}(M,\mathcal{E},\mathcal{F}):=\{u\in H^{1}(M,\mathcal{E}):u|_{\Sigma}\in L^{2}(\Sigma,\mathcal{F})\}.$
By the trace theorem, the restriction map extends uniquely to a bounded linear map
$H^{1}(M,\mathcal{E})\to L^{2}(\Sigma,\mathcal{E}).$
@@
where $\mathcal{F}^{\pm}$ are the eigensubbundles of $\chi$ corresponding to the eigenvalues $\pm 1$.
As usual, we denote by $\nu$ the outward unit normal along $\Sigma$. Let $H$ denote the unnormalized mean curvature. The sign convention is such that the boundary of a Euclidean ball has positive mean curvature. We define an operator $\mathcal{D}^{\Sigma}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
$\mathcal{D}^{\Sigma}=\sum_{j=1}^{n-1}(\gamma(\nu)\otimes\textup{id})(\gamma(e_{j})\otimes\textup{id})\,\nabla_{e_{j}}^{\mathcal{E}}+\tfrac{1}{2}H,$
where $e_{1},\ldots,e_{n-1}$ denotes a local orthonormal frame on $\Sigma$ and $\nabla^{\mathcal{E}}$ denotes the connection on $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ defined above. Moreover, we define an operator $\mathcal{A}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
$\mathcal{A}=\mathcal{D}^{\Sigma}+\tfrac{1}{2}\,\chi\circ\Big(\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j}))\Big).$
Let $\mathcal{S}_{0}$ denote the trivial bundle over $S^{n-1}$ with fiber $S_{0}$, and let $\mathcal{S}_{0}^{*}$ denote the dual bundle over $S^{n-1}$ with fiber $S_{0}^{*}$. With this understood, we may write $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$. By [Baer96]*Proposition 2.2, $\mathcal{D}^{\Sigma}$ can be identified with a twisted Dirac operator acting on sections of
$\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*},$
where the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ carries the flat connection.
Lemma 2.2.
Along $\Sigma$ we have
$\mathcal{D}^{\Sigma}\chi+\chi\mathcal{D}^{\Sigma}=-\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})),$
Proof.
Recall that $\mathcal{D}^{\Sigma}$ anticommutes with $\gamma(\nu)\otimes\textup{id}$. Hence, we compute for a smooth section $s$ of $\mathcal{E}|_{\Sigma}$:
$\mathcal{D}^{\Sigma}\chi s$ $=-\mathcal{D}^{\Sigma}(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))s$
$=(\gamma(\nu)\otimes\textup{id})\mathcal{D}^{\Sigma}(\textup{id}\otimes\gamma_{0}^{*}(N))s$
$=-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})\,\nabla_{e_{j}}^{\mathcal{E}}((\textup{id}\otimes\gamma_{0}^{*}(N))s)+\tfrac{1}{2}H\,(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))s$
$=-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))\,\nabla_{e_{j}}^{\mathcal{E}}s+\tfrac{1}{2}H\,(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))s$
$\quad-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(dN(e_{j})))s$
$=(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))\,\mathcal{D}^{\Sigma}s-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(dN(e_{j})))s$
$=-\chi\mathcal{D}^{\Sigma}s-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(dN(e_{j})))s.\qed$
Lemma 2.3.
Along $\Sigma$ we have $\mathcal{A}\chi+\chi\mathcal{A}=0$.
Proof.
This follows from Lemma 2.2. ∎
The following index computation generalizes the results in [Brendle] and plays a key role in our arguments.
@@
Suppose that $M$ has odd dimension $n\geq 3$. Then the operator